OpenStudy (anonymous):

-2sin x/2 cos x/2 =1-sin x . how do I verify this?

OpenStudy (anonymous):

[cos(x/2) - sin(x/2)]^2 = 1 - sinx cos^2(x/2) - 2cos(x/2)sin(x/2) + sin^2(x/2) cos^2(x/2) + sin^2(x/2) - 2cos(x/2)sin(x/2) 1 - 2cos(x/2)sin(x/2) 1- sin(2x/2) 1- sin(x) <= equal to the right side...

OpenStudy (anonymous):

using the double angle formula: 2sin(u) cos(u) = sin(2u) apply it here............................2 sin(x/2) cos(x/2) = sin[2(x/2)] = sin(x)

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

:)